OpenStudy (anonymous):

NEED ASAP!!! A motorist driving at 25 meters/second decelerates to 15 meters/second. The combined mass of the vehicle and the driver is 1.5 × 10³ kilograms. What is the amount of mechanical work done by the vehicle?

OpenStudy (anonymous):

You can solve this one using the Work-Energy theorem: \[\Delta K = W_{net}\] with \(K\): the kinetic energy, and \(W_{net}\) the net work done in the system. you can find \(\Delta K = K_f-K_i=\frac{1}{2}Mv_f^2-\frac{1}{2}Mv_i^2\) try that!